### v1

# R3-RAG: Learning Step-by-Step Reasoning and Retrieval for LLMs via Reinforcement Learning

Abstract
Retrieval-Augmented Generation (RAG) integrates external knowledge with Large Language Models (LLMs) to enhance factual correctness and mitigate hallucination. However, dense retrievers often become the bottleneck of RAG systems due to their limited parameters compared to LLMs and their inability to perform step-by-step reasoning. While prompt-based iterative RAG attempts to address these limitations, it is constrained by human-designed workflows. To address these limitations, we propose R3-RAG, which uses Reinforcement learning to make the LLM learn how to Reason and Retrieve step by step, thus retrieving comprehensive external knowledge and leading to correct answers. R3-RAG is divided into two stages. We first use cold start to make the model learn the manner of iteratively interleaving reasoning and retrieval. Then we use reinforcement learning to further harness its ability to better explore the external retrieval environment. Specifically, we propose two rewards for R3-RAG: 1) answer correctness for outcome reward, which judges whether the trajectory leads to a correct answer; 2) relevance-based document verification for process reward, encouraging the model to retrieve documents that are relevant to the user question, through which we can let the model learn how to iteratively reason and retrieve relevant documents to get the correct answer. Experimental results show that R3-RAG significantly outperforms baselines and can transfer well to different retrievers. We release R3-RAG at https://github.com/Yuan-Li-FNLP/R3-RAG.
R3-RAG: Learning Step-by-Step Reasoning and Retrieval for LLMs via Reinforcement Learning
Yuan Lithanks: Equal contribution.  Qi Luo11footnotemark: 1  Xiaonan Li11footnotemark: 1  Bufan Li  Qinyuan Cheng Bo WangYining ZhengYuxin WangZhangyue YinXipeng Qiuthanks: Corresponding author. School of Computer Science, Fudan University liyuan24@m.fudan.edu.cnxpqiu@fudan.edu.cn
1 Introduction
Retrieval-Augmented Generation (RAG) has become a prevailing mechanism to address Large Language Models’ (LLMs) hallucination, as shown in Figure 1(a). It typically first retrieves relevant documents and then integrates them to supplement the LLM’s factual knowledge. In this way, RAG can supplement LLMs with external knowledge, mitigating hallucination and leading to better factuality.
(a) Vanilla RAG
(b) Iterative RAG
(c) R3-RAG
Figure 1: Comparison of different RAG approaches: (a) Vanilla RAG: The LLM uses the documents retrieved by the original question to generate the response; (b) Iterative RAG: The LLM interleaves between thinking and invoking the retriever in the fixed, human-designed workflow. (c) R3-RAG: Using reinforcement learning to let the LLM learn how to better reason and retrieve iteratively to get relevant documents and lead to the correct answer.
The response quality of RAG highly depends on the retrieval result  (Li et al., 2024; Yoran et al., 2024). However, the widely used retrievers have become the bottleneck of the RAG system and limited the overall performance. Specifically, dense retrievers usually have significantly fewer parameters and have not been scaled well to the size of LLMs. Additionally, dense retrievers directly generate vectors for input sequences and cannot conduct step-by-step thinking, which makes it challenging to tackle queries that require reasoning. RAG with iterative retrieval (see Figure 1(b)) is proposed to address these limitations. For example, Trivedi et al. (2023) propose to let the LLM interleave retrieval and Chain-of-Thought (CoT) to better retrieve relevant documents. Li et al. (2024) propose to let the LLM iteratively improve the retrieval result until it is sufficient to support answering the user question. However, these methods are limited by human-designed workflows and thus can not fully harness the LLM’s reasoning ability. Besides, LLMs are not trained on how to iteratively invoke retrievers to get comprehensive documents, making the human-designed workflow challenging to synergize the reasoning and retrieval.
To address these limitations, we propose R3-RAG, which uses Reinforcement learning to make the LLM learn how to Reason and Retrieve step by step, thus retrieving comprehensive external knowledge and leading to the correct answer. As shown in Figure 1(c), in R3-RAG, the model can explore better reasoning and retrieval trajectory by the feedback of outcome reward and process reward. Specifically, R3-RAG is divided into two stages. First, we leverage cold start to make the base model learn the manner of iterative reasoning and invoking the retriever. Since cold start makes it difficult for the model to fully explore the external retrieval environment and thus overshadows the model’s ability, we then use RL with outcome rewards to help the model learn how to better reason and retrieve external knowledge step-by-step. To encourage the model to retrieve documents that are relevant to the user question, we further propose relevance-based document verification as process rewards to measure each step’s relevance of reasoning and retrieval. In this way, the model cannot only learn how to obtain the correct answer through outcome rewards, but also reason and retrieve step-by-step through the fine-grained process rewards.
We summarize our contributions as follows:
  • To the best of our knowledge, R3-RAG is the first framework to let the LLM learn how to reason and retrieve step-by-step with outcome rewards and fine-grained process rewards.
  • We conduct comprehensive comparisons with extensive iterative RAG methods and results show that R3-RAG significantly outperforms them, thus indicating its effectiveness.
  • We have released our code and model to facilitate future research at https://github.com/Yuan-Li-FNLP/R3-RAG.
2 Related Work
Retrieval for RAG
The retrieval in RAG can be mainly divided into two categories: 1) Passive Retrieval Guu et al. (2020) implement the Retrieve-Read paradigm where retrievers fetch documents based on input queries before concatenating them with prompts for language model to generate responses, which is followed by many subsequent works  (Lewis et al., 2021; Gao et al., 2023; Li et al., 2023). However, dense retrievers cannot demonstrate reasoning step-by-step, and thus directly using the original question as the query makes it difficult to retrieve comprehensive external information. Iterative Retrieval is proposed to address this limitation, which enables interleaving between thinking and retrieving through decomposed queries. 2) Iterative Retrieval ReAct, proposed by Yao et al. (2023), synchronizes reasoning traces with retrieval actions, while Self-ask, proposed by Press et al. (2023), implements an autonomous question formulation mechanism during the reasoning process. ITER-RETGEN (Shao et al., 2023) introduces cyclic retrieval-generation patterns, FLARE (Jiang et al., 2023) incorporates adaptive retrieval when LLMs generate low-confidence tokens. IRCoT (Trivedi et al., 2023) strategically positions retrieval operations within chain-of-thought reasoning sequences. Compared with these iterative retrieval methods, which are highly limited by the human-designed workflow, R3-RAG can learn how to better explore the external retrieval environment through RL, which significantly addresses the limitation of the human-designed workflow.
Large Reasoning Models
Recently we have seen the emergence of large reasoning models. OpenAI et al. (2024) introduces OpenAI-O1, which demonstrates remarkable step-by-step reasoning and reflection capabilities. Similarly, DeepSeek-AI et al. (2025) develops Deepseek-R1, which is trained by large-scale RL based on outcome rewards. Ma et al. (2025) introduces techniques that enable LLMs to verify their own reasoning steps, leading to more reliable problem-solving capabilities. However, these RL-based methods are highly limited by the base model’s knowledge  (Yue et al., 2025). In this paper, we explore making LLMs learn how to reason and retrieve external information step-by-step, overcoming knowledge limitations of models when addressing complex questions. Guan et al. (2025), Li et al. (2025), Song et al. (2025) and Zheng et al. (2025), concurrently to our work, propose to use RL to improve the iterative RAG. Our method differs in the critical reward design: While they only use the answer correctness as outcome reward, which cannot provide the fine-grained training signal, we additionally propose relevance-based document verification for process reward to better guide the model to iteratively retrieve relevant documents and thus obtain the comprehensive information for the question.
Figure 2: Training Pipeline and Reward Design for R3-RAG
3 Method
RAG struggles to generate high-quality responses when retrievers fail to comprehensively retrieve the required knowledge. However, retrievers are usually weaker than LLMs, thus limiting the LLM’s remarkable ability. To address this limitation, we propose R3-RAG, which uses Reinforcement learning to make the LLM learn how to Reason and Retrieve step by step, thus retrieving the comprehensive external knowledge and generating the factual answer. Specifically, R3-RAG is divided into two stages, shown in Figure 2. First, we leverage cold start (DeepSeek-AI et al., 2025) to make the base model learn the manner of iterative reasoning and invoking retriever. Since cold start makes the model difficult to fully explore the external retrieval environment, we then use RL to further harness its ability to better search the relevant external knowledge. Specifically, we propose two rewards for R3-RAG: 1) answer correctness for outcome reward, which judges whether the trajectory leads to correct answers by ground truth; 2) relevance-based document verification for proess reward, encouraging the model to retrieve documents that are relevant to the user question. In this section, we first introduce the trajectory of R3-RAG and then the cold start and RL training.
3.1 Trajectory Definition
In this section, we introduce the trajectory definition of R3-RAG. Given a multi-hop question-answering (QA) dataset $\mathcal{Q}=\{(q_{i},a_{i}^{*})\}_{i=1}^{n}$ where $q_{i}$ is a question and $a_{i}^{*}$ is its golden answer, we define a trajectory $T_{i}$ for each question $q_{i}$ as a sequence of reasoning and retrieval steps $T_{i}=\{s_{i}^{1},s_{i}^{2},\ldots,s_{i}^{|T_{i}|}\}$. Each step $s_{i}^{j}$ in trajectory $T_{i}$ is as follows:
$s_{i}^{j}=\begin{cases}(\xi_{i}^{j},\,\psi_{i}^{j},\,d_{i}^{j})&j<\lvert T_{i}\rvert\\[4.0pt] (\xi_{i}^{j},\,a_{i})&j=\lvert T_{i}\rvert\end{cases},$ (1)
where the model iteratively reasons and invokes retrievers until it thinks the documents are enough for generating the answer at step $\lvert T_{i}\rvert$. $\boldsymbol{\xi_{i}^{j}}$ denotes the Reasoning process at step $j$, which analyzes all information, e.g., analysis and documents accumulated up to the step $j-1$ and plans the step $j$’s action; $\boldsymbol{\psi_{i}^{j}}$ is the Retrieval Query produced from this reasoning; $\boldsymbol{d_{i}^{j}}$ comprises the Documents returned by the retriever for that query; and $\boldsymbol{a_{i}}$ is the final Answer to the original question.
During the reasoning process $\boldsymbol{\xi_{i}^{j}}$, the model can step-by-step perform analysis of the information gathered in previous steps. If the existing documents are enough to support the question, the model will output the answer $\boldsymbol{a_{i}}$; otherwise, it determines what additional information is needed and invokes the retriever with a new query $\boldsymbol{\psi_{i}^{j}}$.
3.2 Cold Start
Since base models cannot yet interleave step-by-step reasoning with retrieval, we first employ a cold start stage for R3-RAG. We show the detailed cold start data synthesis in Algorithm 1. For each question, we build its response which follows the defined trajectory iteratively. In each iteration, we first let a strong LLM generate the reasoning step and corresponding action, divided into two cases: 1) the reasoning process and retrieval query, or 2) the reasoning process and final answer. If the LLM gives the final answer, which indicates the LLM thinks it has sufficient information to answer the question, we will finish the question’s response generation. After the response generation, each question has a corresponding trajectory of iterative reasoning and retrieving processes, accompanied by its predicted answer. Additionally, we perform rejective sampling (DeepSeek-AI et al., 2025) based on the question’s gold answer to filter out trajectories that lead to the wrong answer, which further improves the quality of cold-start data. Additional data generation details are available in Appendix A. By performing SFT with the cold-start dataset, we enable the model to learn the manner of iteratively reasoning and retrieval, which can help the model better sample positive trajectories and thus obtain higher RL training signal. Further training details are provided in Appendix B.
Algorithm 1 Cold-Start Data Generation
0:  Dataset $\mathcal{Q}_{CS}=\{(q_{i},a_{i}^{*})\}_{i=1}^{n}$, Retriever $R$, Language Model $M$
0:  Trajectory-Augmented Dataset $\mathcal{D}_{CS}$
1:  $\mathtt{MAX\_ITERATION}\leftarrow 5$$\mathcal{D}_{CS}\leftarrow\emptyset$
2:  for each $(q_{i},a_{i}^{*})\in\mathcal{Q}_{CS}$ do
3:     $T_{i},\mathtt{flag}\leftarrow[\,],\mathtt{False}$
4:     for turn $\in[1,\mathtt{MAX\_ITERATION}]$ do
5:        $\mathtt{step}\leftarrow\mathtt{Parse}(M(\mathtt{Prompt}(q_{i},T_{i})))$
6:        if $\mathtt{{}^{\prime}answer^{\prime}}\in\mathtt{step}$ then
7:           $T_{i}.\mathtt{append}(\mathtt{step})$
8:           break
9:        else if $\mathtt{{}^{\prime}query^{\prime}}\in\mathtt{step}$ then
10:           $\mathtt{step}[\mathtt{{}^{\prime}doc^{\prime}}]\leftarrow R(\mathtt{step}[\mathtt{{}^{\prime}query^{\prime}}])$
11:           $T_{i}.\mathtt{append}(\mathtt{step})$
12:           if $\mathtt{turn}==\mathtt{MAX\_ITERATION}$ then
13:              break {Fail to generate the answer}
14:           end if
15:        else
16:           break {The format of $T_{i}$ is invalid}
17:        end if
18:     end for
19:     $\mathcal{D}_{CS}\leftarrow\mathcal{D}_{CS}\cup\{(q_{i},a_{i}^{*},T_{i})\}$
20:  end for
21:  return  $\mathcal{D}_{CS}$
3.3 Reinforcement Learning
Although the cold start can let the model learn the manner of iterative reasoning and retrieval, it is hard to make the model fully explore the external retrieval environment and thus makes the model limited by the cold-start data. To address these challenges, we propose to use RL to further harness the model’s own abilities to better reason and retrieve external information. Specifically, we propose two rewards for R3-RAG: 1) answer correctness for outcom rewards, which judge whether the trajectory leads to correct answers by ground truth; 2) relevance-based document verification for process reward, encouraging the model to retrieve documents that are highly relevant to user questions. In this section, we first introduce the reward design and then the training objective.
Answer Correctness
The high-quality outcome reward is critical for the model to generate the correct trajectory  (DeepSeek-AI et al., 2025). For question-answering, string matching is usually used to judge whether the prediction is correct according to the gold answer. However, since the same entity can have different variants, the match-based judgment can be inaccurate, which introduces extra noise to the training signal. In this paper, we propose to combine the match-based judgment and model-based judgment as follows:
$\mathrm{Acc}(a)=\begin{cases}1&\text{if }\mathrm{Acc}_{match}(a)\operatorname{or}\mathrm{Acc}_{model}(a)\\ 0&otherwise\end{cases},$ (2)
where $a$ is the predicted answer, and $\mathrm{Acc}_{match}$ and $\mathrm{Acc}_{model}$ are match-based and model-based answer judgement.
Document Relevance
Using only outcome reward is challenging because it provides limited fine-grained training signal and results in sparse rewards, which degrades the effect of RL. We further propose relevance-based document verification to give fine-grained guidance to the process of iterative retrieval. In each retrieval step, we calculate the document relevance as:
$Rel(d_{i}^{j})=\mathrm{LLM}(I_{relevance},q_{i},d_{i}^{j}),$ (3)
where $I_{relevance}$ is the instruction for the LLM to judge the relevance between question $q_{i}$ and documents $d_{i}^{j}$, with scores ranging from $0,0.1,0.2,\dots$ to $1$. Since $q_{i}$ and $d_{i}^{j}$ are concatenated and then sent to the LLM, the relevance judgment can take advantage of full token-level interaction and thus guide the model to better retrieve relevant information. Additionally, since retrieving relevant documents in one retrieval step is easier than getting the correct answer in the final step, this relevance-based reward can provide more fine-grained guidance and make it easier for the model to learn how to retrieve comprehensive information.
Format Correctness
We also implement a format reward $r_{format}(s_{i}^{j})$ to encourage the model to generate the retrieval invoking in the required format.
Overall Reward
The overall reward is the sum of the answer correctness reward, document relevance reward, and format reward. In this way, we can let the model learn how to iteratively invoke retrievers and obtain external information to achieve the final correct answer. To enhance the effect of answer correctness based reward and prevent the training progress from being distracted by other rewards, we use the following factor to adjust the process reward throughout the trajectory:
$r^{*}_{\text{process}}(s_{i}^{j})=r_{\text{process}}(s_{i}^{j})\cdot\lambda_{T}(T_{i}),$ (4)
where $\lambda_{T}(T_{i})$ is determined based on the trajectory outcome. For trajectories with correct answers, $\lambda_{T}(T_{i})$ is positive to reinforce successful reasoning; for incorrect answers or failures to produce answers due to overthinking, it is negative to discourage these behaviors. For trajectories with format errors, $\lambda_{T}(T_{i})$ equals 1 to isolate format penalties to specific steps.
3.3.1 Training Objective
We apply the Proximal Policy Optimization (PPO) on the cold-start model. Guided by rewards from both the reasoning process and the final outcome, we update the model weights according to the following loss function:
$\mathcal{L}_{RL}=$ $\mathbb{E}{(x,y)\sim\mathcal{D}_{RL\_train}}\Big{[}\min\Big{(}\rho(\theta)A(x,y),$
$\text{clip}(\rho(\theta),1-\epsilon,1+\epsilon)A(x,y)\Big{)}\Big{]}$ (5)
$+\beta\mathcal{L}_{KL},$
where $\rho(\theta)=\frac{P(y|x;\theta)}{P(y|x;\theta_{old})}$ represents the probability ratio between the current policy and the old policy, $A(x,y)$ denotes the advantage function estimated using Generalized Advantage Estimation (GAE), $\epsilon$ is the clipping parameter that constrains policy updates, $\mathcal{L}_{KL}$ represents the Kullback-Leibler divergence between the current policy and the reference model to prevent excessive deviation, and $\beta$ controls the strength of the KL penalty (Schulman et al., 2017). Details are provided in Appendix D.
4 Experiments
HotpotQA 2WikiMultiHopQA MuSiQue Average
Methods Retriever ACC F1 EM ACC F1 EM ACC F1 EM ACC
Llama-3.1-8B
CoT - 39.2 38.3 27.6 28.8 34.5 25.3 14.0 17.2 8.0 27.3
RAG with CoT E5 53.3 50.8 38.6 32.9 40.2 31.1 16.3 18.8 10.3 34.2
ReAct(Tool Call) E5 30.8 25.8 21.2 17.8 14.5 12.9 8.6 6.5 4.8 19.1
Flare E5 24.6 20.9 17.8 11.9 11.4 10.9 3.5 2.8 2.3 13.3
Self-ask E5 39.1 34.3 29.5 29.0 27.5 26.2 14.4 23.5 13.5 27.5
ITER-RETGEN E5 45.7 40.0 34.3 19.4 18.1 17.0 11.1 8.9 7.9 25.4
IRCoT E5 52.8 46.0 39.3 40.6 37.5 35.1 16.7 13.6 12.0 36.7
Auto-RAG BM25 - 36.1 25.8 - 30.1 23.0 - - - -
DeepRAG BM25 - 51.5 40.7 - 53.3 48.1 - - - -
R3-RAG-CS E5 60.6 55.5 43.6 53.7 53.4 46.5 29.6 29.4 19.4 48.0
R3-RAG E5 64.4 58.8 45.6 61.0 60.9 52.9 32.2 32.7 21.1 52.6
R3-RAG BM25 62.5 57.6 44.4 58.0 58.6 50.6 26.4 27.7 17.2 49.0
R3-RAG BGE 65.3 60.0 46.6 62.1 61.8 53.7 33.8 32.8 21.2 53.8
Qwen2.5-7B
CoT - 34.0 34.0 23.0 31.1 36.9 27.4 12.7 17.0 6.7 25.9
RAG with CoT E5 52.4 49.4 37.6 33.5 39.1 30.1 16.9 18.8 9.9 34.3
ReAct(Tool Call) E5 27.9 23.1 21.1 38.3 31.7 26.1 11.7 8.0 5.8 26.0
Flare E5 28.9 24.8 21.2 29.1 28.3 27.8 6.5 4.8 3.6 21.5
Self-ask E5 37.1 33.2 28.7 31.5 29.8 28.2 13.3 10.7 8.6 27.3
ITER-RETGEN E5 48.4 40.2 35.8 32.8 31.4 30.2 12.3 10.1 8.5 31.2
IRCoT E5 48.4 41.1 35.7 35.8 33.5 31.1 13.5 11.2 9.4 32.6
DeepRAG BM25 - 41.1 32.1 - 44.9 40.4 - - - -
ReSearch-7B BGE 60.3 - 40.6 50.1 - 44.7 32.2 - 21.7 47.5
ReSearch-7B-Instruct BGE 63.6 - 43.5 54.2 - 47.6 33.4 - 22.3 50.4
R3-RAG-CS E5 63.3 57.6 45.2 53.1 53.5 46.1 31.3 31.9 21.9 49.2
R3-RAG E5 65.5 59.7 46.4 62.3 62.7 54.2 33.6 34.0 21.4 53.8
R3-RAG BM25 63.8 58.2 44.9 59.6 61.1 52.8 29.2 30.0 17.6 50.8
R3-RAG BGE 66.4 60.6 46.8 63.0 63.4 55.2 34.8 34.3 21.7 54.8
Table 1: Performance comparison of different models and retrieval methods across multi-hop QA datasets. Our R3-RAG method consistently outperforms all baselines across accuracy (ACC), F1, and exact match (EM) metrics regardless of the base model or retriever used. Bold values indicate best performance in each column and model group. Missing values (-) indicate results not reported in original papers.
4.1 Datasets and Metrics
We follow Song et al. (2025) to conduct comparison on three multi-hop QA datasets: HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022), using their full dev sets. We evaluate using three complementary metrics: Exact Match (EM) measures strict string match with golden answers; F1 Score quantifies partial matching through token overlap between predicted and golden answers; and Accuracy (Acc) is based on LLM judgment of semantic correctness. For EM and F1, we first extract the short answer from the model’s response and then compare it with the gold answer.
4.2 Baselines
We compare against several competitive baselines: (1) CoT: the LLMs directly answer the question with Chain-of-Thought; (2) RAG with CoT: the LLMs answer the question based on the documents retrieved by original questions as queries; (3) Iterative retrieval methods: models are prompted to invoke retrieval tools during reasoning, including ReAct Yao et al. (2023), Flare Jiang et al. (2023), Self-ask Press et al. (2023), ITER-RETGEN Shao et al. (2023), and IRCoT Trivedi et al. (2023); and (4) Fine-tuned methods: including Auto-RAG Yu et al. (2024), DeepRAG Guan et al. (2025), and ReSearch Chen et al. (2025).
4.3 Implementation Details
For the cold-start stage, we create 51,254 solution trajectories from the training sets of HotpotQA Yang et al. (2018), 2WikiMultiHopQA Ho et al. (2020), and MuSiQue Trivedi et al. (2022). For the reinforcement learning stage, we sample 8,192 instances from the remaining HotpotQA training data. We employ E5-base-v2 (Wang et al., 2022) as our default retriever. During evaluation, the models receive complete retrieval documents, limited to the top-5 per iteration with a maximum of 5 iterations. Unlike prior work  (Zheng et al., 2025), which extracts information from retrieval documents to reduce context length, our approach directly incorporates complete retrieval documents. All methods are evaluated using the same metrics and hyper-parameters. Further implementation details are provided in Appendix E.
4.4 Main Results
We show the evaluation results of R3-RAG using Llama-3.1-8B and Qwen2.5-7B in Table 1. We see that R3-RAG significantly outperforms baselines on three datasets across different metrics, which shows its best overall reasoning and retrieval performance. Specifically, R3-RAG outperforms the strongest iterative RAG workflow method, IRCOT, by around 15 points on average, which shows that R3-RAG can help the model better learn how to iteratively reason and retrieve step-by-step and thus get more accurate answers. Additionally, R3-RAG also outperforms ReSearch (Chen et al., 2025), which uses RL with outcome-based reward to augment the model’s iterative retrieval ability. This shows the advantages of R3-RAG’s reward design, which not only teaches the model how to get the final answers, but also guides the model to retrieve documents relevant to the question. Meanwhile, R3-RAG achieves significant improvement based on both Qwen2.5-7B and Llama-3.1-8B, and this shows the robustness of R3-RAG. Although R3-RAG’s RL is only trained on the HotpotQA dataset, R3-RAG outperforms the cold-start model on all three datasets, which shows the proposed RL algorithm can generalize well across various datasets and further demonstrates R3-RAG’s transferability for different downstream scenarios.
5 Analysis
5.1 Effect of Different Reward
Methods 2Wiki MuSiQue Avg
R3-RAG 61.0 32.2 46.6
   - Process Reward 59.0 31.4 45.2
       - Outcome Reward 53.7 29.6 41.6
Table 2: Effect of Different Reward on R3-RAG.
Table 2 presents the comparative performance of different reinforcement learning configurations in R3-RAG. Our full model employs both process and outcome rewards. When we remove the process reward component from our complete R3-RAG framework, we observe a performance decline of 1.4% on average (from 46.6% to 45.2%), which shows that fine-grained process rewards effectively guide the model to retrieve more relevant documents at each reasoning step, leading to improved overall performance. Further removing the outcome reward component (thus eliminating all reinforcement learning signals) results in an additional performance drop of 3.6% on average (from 45.2% to 41.6%). These results demonstrate the effectiveness of both reward components. Outcome rewards guide the model to optimize for correct final answers, while process rewards provide granular feedback on the quality of each retrieval operation. By jointly optimizing for both the ultimate goal (correct answers) and the intermediate steps (relevant document retrieval), our approach achieves a more effective reasoning-retrieval strategy.
Figure 3: Impact of the maximum number of iteration steps on HotpotQA and 2WikiMultiHopQA. Results are shown for both Qwen and Llama backbone models. All models are trained with up to five reasoning steps.
5.2 Effect of Maximum Iteration Steps
In this section, we analyze how the number of iteration steps affects model performance. As shown in Figure 3, R3-RAG demonstrates steady improvements in answer accuracy as the number of steps increases from 1 to 10. Notably, R3-RAG continues to benefit from additional reasoning steps at inference time, even beyond the maximum number used during training (five steps). This suggests that our framework enables models to generalize their reasoning procedures effectively to longer inference trajectories. Furthermore, we observe that R3-RAG consistently outperforms R3-RAG-CS when evaluated with the same maximum number of steps, highlighting the effectiveness of our RL training approach in enabling models to more accurately answer questions while utilizing an equivalent number of reasoning iterations.
5.3 Impact of Retrieval Top-k Parameter
Figure 4: R3-RAG performance across different retrieval top-k.
We analyze how the number of documents retrieved at each step affects our method’s performance. Figure 4 demonstrates that increasing top-k from 1 to 5 significantly improves performance by approximately 8%, while adding more documents beyond 5 yields minimal gains (less than 1%). Additionally, we observe that our RL approach substantially enhances answer accuracy across all top-k configurations, demonstrating the robustness of our RL method regardless of retrieval volume.
Methods Retriever Hotpot 2Wiki Avg
Llama-3.1-8B
CoT with RAG BM25 52.3 31.5 41.9
E5 53.3 32.9 43.1
BGE 55.8 33.9 44.8
ReAct BM25 33.6 18.0 25.8
E5 30.8 17.8 24.3
BGE 29.8 16.1 22.9
IRCoT BM25 52.8 41.4 47.1
E5 52.8 40.6 46.7
BGE 43.3 20.8 32.0
R3-RAG BM25 62.5 58.0 60.2
E5 64.4 61.0 62.7
BGE 65.3 62.1 63.7
Qwen2.5-7B
CoT with RAG BM25 49.9 29.8 39.9
E5 52.4 33.5 42.9
BGE 55.1 35.3 45.2
ReAct BM25 43.4 28.7 36.1
E5 27.9 38.3 33.1
BGE 37.7 27.4 32.6
IRCoT BM25 48.5 33.8 41.2
E5 48.4 35.8 42.1
BGE 40.5 20.1 30.3
R3-RAG BM25 63.8 59.6 61.7
E5 65.5 62.3 63.9
BGE 66.4 63.0 64.7
Table 3: Performance comparison of different retrieval methods (BM25, E5, BGE) across models on HotpotQA and 2WikiMultiHopQA datasets. Results show accuracy (ACC) metric only.
5.4 Transferability over Different Retrievers
We analyze the impact of different retrievers on our method compared with several baselines and show the results in Table 3. Despite being trained exclusively with the E5 retriever, R3-RAG maintains the performance and consistently outperforms all baselines when using different retrievers, demonstrating the transferability of R3-RAG’s learned reasoning and retrieval strategies across external retrieval environment. Meanwhile, R3-RAG’s performance variance across three different retrievers is lower than 3%, which further shows R3-RAG’s robustness. In contrast, baseline methods show significant fluctuation. For example, IRCoT’s accuracy on the 2Wiki dataset plummets from 35.8% when using E5 to just 20.1% with BGE, revealing the human-designed workflow’s sensitivity to different external retrieval environment. Additional results are provided in Appendix F.
5.5 Case Study
Figure 5: Case Study: R3-RAG can reason and retrieve step by step, and adaptively modify the query when it meets retrieval failure.
We conduct a case study on a multi-hop question, as shown in Figure 5, which asks for the film, The Model Couple or Thendral Veesum, whose director was born earlier. RAG + CoT fails to retrieve the comprehensive information. Specifically, it only retrieves who are these movies’ directors and thus fails to generate the answer. In contrast, R3-RAG can decompose the whole question into the following chain of retrieval queries: 1) the director of the film The Model Couple -> William Klein; 2) the director of the film Thendral Veesum -> B. S. Ranga; 3) the birth day of William Klein 4) the birth day of B. S. Ranga. By iterative reasoning and retrieval, R3-RAG can retrieve relevant information progressively and thus get the comprehensive information of the given question. Additionally, R3-RAG can adaptively decompose the previous query when it fails to retrieve relevant information. Specifically, then it tries to retrieve the documents which contain both directors’ birth day and fails, it can further decompose this query into two separate queries and then successfully gets the corresponding information, which shows R3-RAG’s potential when it meets complicated queries.
6 Conclusion
We propose R3-RAG, a framework that leverages reinforcement learning to teach LLMs how to reason and retrieve step by step, optimizing the interleaved reasoning and retrieval process with both outcome rewards and fine-grained process rewards through relevance-based document verification. Experimental results demonstrate that R3-RAG significantly outperforms extensive iterative RAG methods across multiple benchmarks. Further analyses show the effectiveness of R3-RAG’s reward design and its transferability across different retrievers.
Limitations
Despite the effectiveness of R3-RAG, our approach has two main limitations. First, our experiments rely on publicly available academic datasets, which may lack sufficient diversity and realism, potentially limiting the model’s robustness and generalization to real-world user queries. Future research could address this by curating more diverse and representative datasets. Second, our cold-start data quality is constrained by the reasoning capabilities of the foundation model (DeepSeek-v3). Due to its limitations, we could not reliably generate correct reasoning trajectories for certain challenging multi-hop questions, even after repeated sampling. Using stronger models, such as DeepSeek-R1, could further enhance cold-start data quality in future work.
Ethics Statement
Our research complies with ethical standards in NLP research. We utilize publicly available datasets (HotpotQA, 2WikiMultiHopQA, and MuSiQue) and models (Llama-3.1-8B, Qwen2.5-7B) with appropriate licenses for academic research. Our derivative models maintain compatibility with their original licenses. Detailed information about datasets, models, frameworks, and licensing can be found in Appendix H.
Regarding data privacy and content safety, we conducted a thorough review of the datasets to ensure they do not contain personally identifying information or offensive content. The datasets consist of questions and answers derived from Wikipedia, which undergoes content moderation. The questions focus on factual knowledge and do not target sensitive personal information. No additional data collection involving human subjects was conducted for this research, eliminating risks related to personal data exposure. Furthermore, we did not implement any mechanisms that could potentially generate harmful or offensive content. The training objectives were focused on improving factual accuracy and reasoning capabilities for answering multi-hop questions based on retrieved information.
Appendix A Data Generation Details of Cold Start
The complete prompt templates used in Algorithm 1 are shown in Figure 6 and Figure 7. The process consists of two main parts. In the first step as illustrated in Figure 6, the model analyzes the problem comprehensively, planning the solution trajectory and deciding whether to generate a retrieval query or provide a final answer based on information sufficiency. For subsequent reasoning steps presented in Figure 7, the prompt incorporates previous reasoning, queries, and retrieved documents, guiding the model to evaluate retrieved information, identify knowledge gaps, and determine whether additional retrieval is needed or if sufficient information has been gathered to answer the question.
After obtaining an initial dataset through Algorithm 1, to further enhance data quality, we vary the model’s sampling temperature to produce multiple samples for each problem. We then evaluate these different solution paths using Best-of-N (BON) selection, ultimately choosing the highest quality trajectory that correctly arrives at the correct answer, adheres to the proper format, and uses the most concise reasoning steps possible. Through this process, we obtain the dataset $\mathcal{D}_{CS}$ finally used for cold start.
Algorithm 2 Training Data Generation
0:  Trajectory-augmented dataset $\mathcal{D}_{CS}$
0:  Imitation learning training dataset $\mathcal{D}_{CS\_train}$
1:  $\mathcal{D}_{CS\_train}\leftarrow\emptyset$
2:  for each $(q_{i},a_{i}^{*},T_{i})\in\mathcal{D}_{CS}$ do
3:     $\mathtt{context}\leftarrow q_{i}$
4:     $\mathtt{history}\leftarrow\emptyset$
5:     for $j=1$ to $|T_{i}|$ do
6:        $\mathtt{input}\leftarrow\mathtt{context}\cup\mathtt{history}$
7:        if $j<|T_{i}|$ then
8:           $(\xi_{i}^{j},\,\psi_{i}^{j},\,d_{i}^{j})\leftarrow s_{i}^{j}$
9:           $\mathtt{output}\leftarrow\mathtt{Format}(\xi_{i}^{j},\,\psi_{i}^{j},\,d_{i}^{j})$
10:        else
11:           $(\xi_{i}^{j},\,a_{i})\leftarrow s_{i}^{j}$
12:           $\mathtt{output}\leftarrow\mathtt{Format}(\xi_{i}^{j},\,a_{i})$
13:        end if
14:        $\mathtt{history}\leftarrow\mathtt{history}\cup\{\mathtt{output}\}$
15:        $\mathtt{pair}\leftarrow(\mathtt{input},\mathtt{output})$
16:        $\mathcal{D}_{CS\_train}\leftarrow\mathcal{D}_{CS\_train}\cup\{\mathtt{pair}\}$
17:     end for
18:  end for
19:  return  $\mathcal{D}_{CS\_train}$
For each trajectory in $\mathcal{D}_{CS}$, we generate multiple training examples where each example consists of the original question concatenated with all previous reasoning steps as input, and the current reasoning step as output. This approach naturally addresses the masking requirement of retrieval documents, as retrieved documents are included in the input context during training, and the input context is masked during the training process. We define this processed training dataset as $\mathcal{D}_{CS\_train}=\{(\mathtt{input}_{k},\mathtt{output}_{k})\}_{k=1}^{m}$, where $m=\sum_{i=1}^{|\mathcal{Q}^{\prime}_{CS}|}|T_{i}|$. Algorithm 2 illustrates our approach to converting the trajectory-augmented dataset $\mathcal{D}_{CS}$ into this training dataset $\mathcal{D}_{CS\_train}$. For each data $(q_{i},a_{i}^{*},T_{i})$ in $\mathcal{D}_{CS}$, we format the structured data $s_{i}^{j}$ in each step of $T_{i}$ into strings that conform to our required format. For a trajectory $T_{i}$, we iterate through each step, using all steps preceding the current one as the input for the training dataset, and the content of the current step as the output, ultimately creating $|T_{i}|$ training examples.
Appendix B Training Details of Cold Start
Based on the training dataset $\mathcal{D}_{CS\_train}$ generated from Algorithm 2, we implement Supervised Fine-Tuning (SFT) to train our model. Applying the maximum likelihood estimation principle, the loss function can be formulated as
$\mathcal{L}_{CS}=-\sum_{(x,y)\in\mathcal{D}_{CS\_train}}\log P(y|x;\theta),$ (6)
where $x$ is equal to $\mathtt{input}$ representing the combination of the question and all previous reasoning steps of the question, $y$ is equal to $\mathtt{output}$ representing the current step output of the question, and $\theta$ denotes the model parameters. When expanded across all questions in our trajectory-augmented dataset $\mathcal{D}_{CS}$, this loss function becomes
$\mathcal{L}_{IR}=-\sum_{i=1}^{|\mathcal{Q}^{\prime}_{CS}|}\sum_{j=1}^{|T_{i}|}\log P(s_{i}^{j}|h_{i}^{j};\theta),$ (7)
where $h_{i}^{j}=(q_{i},s_{i}^{1},\ldots,s_{i}^{j-1})$ represents the history of steps preceding step $j$ for question $i$, with the convention that $s_{i}^{0}=\emptyset$ for all $i$. The trained model is called as $\mathcal{R}^{3}$-RAG-CS.
Appendix C Online Data Sampling of RL
Unlike the training of cold start or offline RLHF where data can be prepared in advance, PPO requires online sampling during the training process, with the dataset divided into multiple batches for sampling and training.
We use multi-hop question-answering datasets $\mathcal{Q}_{RL}$ that are distinct from the cold-start datasets ($\mathcal{Q}_{RL}\cap\mathcal{Q}_{IL}=\emptyset$) for reinforcement learning, with online sampling logic consistent with the cold-start phase. We denote the sampled dataset as $\mathcal{D}_{RL}=\{(q_{i},a_{i}^{*},T_{i})\mid(q_{i},a_{i}^{*})\}$ and the final training dataset as $\mathcal{D}_{RL\_train}=\{(\mathtt{input}_{k},\mathtt{output}_{k})\}_{k=1}^{m^{\prime}}$, where $m^{\prime}=\sum_{i=1}^{|\mathcal{Q}_{RL}|}|T_{i}|$, with each pair consisting of the question with previous reasoning steps as input and the current reasoning step as output.
A key distinction from the cold-start phase is that during reinforcement learning, we preserve all trajectories regardless of correctness or completeness. Consequently, the final step in any trajectory $T_{i}$ from $\mathcal{D}_{RL}$ may lack an answer or generate an incorrect answer.
Appendix D Generalized Advantage Estimation and KL Divergence Definition
The advantage function $A(x,y)$ in reinforcement learning is used to guide policy improvement. We estimate $A(x,y)$ using Generalized Advantage Estimation (GAE). In GAE, the advantage $A(x,y)$ is defined as follows:
$A(x,y)=\sum_{t=0}^{T-1}(\gamma\lambda)^{t}\delta_{t},$ (8)
where $\delta_{t}$ is the Temporal Difference (TD) error at time step $t$, which is $\delta_{t}=r_{t}+\gamma V(y)-V(x)$. Specifically, $x$ represents the current state $s_{t}$, $y$ represents the next state $s_{t+1}$, $r_{t}$ is the reward received at time step $t$, $\gamma$ is the discount factor, $V(x)$ and $V(y)$ are the value functions for states $x$ and $y$, respectively, and $\lambda$ is a hyperparameter that controls the trade-off between bias and variance in the advantage estimation process. The summation in the definition of $A(x,y)$ captures the idea of taking into account not just the immediate rewards but also the future rewards, discounted over time, to provide a more accurate estimation of the advantage.
The term $\mathcal{L}_{KL}$ represents a KL-divergence penalty to prevent the policy from deviating too much from the initial model:
$\mathcal{L}_{KL}=D_{KL}[P(y|x;\theta_{init})\parallel P(y|x;\theta)],$ (9)
where $\beta$ controls the strength of this regularization. The details of the process reward model will be provided in the next section.
Appendix E Implementation Details
We use Llama-3.1-8B Grattafiori et al. (2024) and Qwen2.5-7B Team (2024) models for training and evaluating. For the stage of the cold start, we prompt DeepSeek-V3 DeepSeek-AI (2024) models to obtain step-by-step solution trajectories that can solve problems correctly with a specific format. We sampled from 15,000 examples in the 2WikiMultiHopQA training set, 19,937 examples in the MuSiQue training set, and the first 30,000 examples in HotpotQA, resulting in 10,843, 14,145, and 26,266 solution trajectories respectively. Then we further cleaned these data to ensure they conform to our required format specifications. We obtained a total of 51,254 trajectories for cold start, which can arrive at correct answers but do not necessarily follow optimal solution paths. We use these lower-quality data to train our base model through the LLaMA-Factory framework Zheng et al. (2024), resulting in the first version of our method, which we call R3-RAG-CS. During training, we apply full-parameter fine-tuning with DeepSpeed zero-3 optimization and a maximum token length of 4,096. For Llama and Qwen models, we set learning rates at 1.0e-5 and 7.0e-6, with batch sizes of 8 and 16 respectively.
For the stage of RL, since we have fully utilized the 2WikiMultiHopQA and MuSiQue training datasets, we extract 8,192 samples from the remaining 60,000+ HotpotQA training data for reinforcement learning. We modified the sampling code in the OpenRLHF framework Hu et al. (2024) to enable step-by-step sampling and tool calling according to our requirements. Then we use the modified OpenRLHF to complete PPO reinforcement learning training, resulting in our R3-RAG models. Actor learning rates are set at 5e-7 for both models, critic learning rates at 9e-6, and the training batch size is uniformly set at 64. The context length after multiple inference iterations and model calls may lead to CUDA out of memory errors. To avoid frequent errors, we limit each retrieved document to 512 characters during reinforcement learning training. we set $\gamma_{fail}=-0.1$, $\gamma_{incorrect}=-0.2$, and $\gamma_{correct}=0.6$. And we use Qwen2.5-14B-Instruct Team (2024) to assign scores of document relevance.
We use the wikipedia-nq-corpusfrom Tevatron (wikipedia2018) with the FlashRAG Jin et al. (2024) framework to build the retriever. If the retriever is not specified, we default to using the E5-base-v2 modelWang et al. (2022) as the retriever. Additionally, for fair testing, we also provide results using BGE-large-en-v1.5Xiao et al. (2023) and BM25Lù (2024). When using the E5 and BGE models, We use the faiss library Johnson et al. (2019) for GPU-accelerated retrieval.
In evaluation, although we know that excessive context length can degrade model reasoning performance, we directly input the retrieved content into the model without extracting key information as other methods do. This better represents the model’s capabilities in real-world environments, where useful information is often hidden among large amounts of irrelevant information. We set the retriever’s top-k to 5 and limit the maximum number of iterations to 5 during evaluation. Specifically, for the Naive RAG method, since each question only retrieves once, we set the number of top-k to 16, as our proposed method uses fewer than 16 documents on average. If the model fails to answer a question within this number of steps, we consider it unable to answer the question and count it as incorrect. We implemented Naive Generation, Naive RAG, and react methods, prompting the models to generate answers step by step. For baselines including Flare, Self-ask, ITER-RETGEN, and IRCoT, we used the FlashRAG implementation. All methods first generate answers, which are then evaluated using a unified evaluation script. Evaluation data for Auto-RAG and DeepRAG are cited from the DeepRAGGuan et al. (2025), while ReSearch-Qwen-7B-Instruct results are cited from its paperChen et al. (2025). We ensured consistent evaluation parameters across all methods. And we use Qwen2.5-72B-Instruct Team (2024) to judge the correctness for the accuracy in our metrics.
We conducted all our model training experiments twice to ensure the training process is reproducible. During evaluation, we observe that the model’s output results were stable, therefore we only performed a single trial. We conduct our experiments using NVIDIA H100 GPUs.
Appendix F Details of Analysis for Retriever Quality
HotpotQA 2WikiMultiHopQA MuSiQue Avg
Methods Retriever ACC F1 EM ACC F1 EM ACC F1 EM ACC
Llama-3.1-8B
RAG with CoT BM25 52.3 49.9 38.1 31.5 39.4 30.5 13.3 16.3 8.0 32.4
E5 53.3 50.8 38.6 32.9 40.2 31.1 16.3 18.8 10.3 34.2
BGE 55.8 52.8 40.5 33.9 41.1 32.4 16.8 19.5 10.3 35.5
ReAct(Tool Call) BM25 33.6 28.3 23.3 18.0 15.4 13.6 8.4 5.8 4.3 20.0
E5 30.8 25.8 21.2 17.8 14.5 12.9 8.6 6.5 4.8 19.1
BGE 29.8 25.2 20.6 16.1 13.2 11.6 8.2 6.0 4.6 18.0
IRCoT BM25 52.8 46.4 39.5 41.4 38.6 36.2 17.5 14.0 12.0 37.2
E5 52.8 46.0 39.3 40.6 37.5 35.1 16.7 13.6 12.0 36.7
BGE 43.3 37.8 32.4 20.8 18.9 17.9 12.8 10.3 9.0 25.6
R3-RAG-CS BM25 57.8 53.3 41.6 48.0 49.1 42.3 22.6 23.0 15.4 42.8
E5 60.6 55.5 43.6 53.7 53.4 46.5 29.6 29.4 19.4 48.0
BGE 61.7 56.8 44.5 54.6 54.4 47.4 28.8 28.9 19.4 48.4
R3-RAG BM25 62.5 57.6 44.4 58.0 58.6 50.6 26.4 27.7 17.2 49.0
E5 64.4 58.8 45.6 61.0 60.9 52.9 32.2 32.7 21.1 52.5
BGE 65.3 60.0 46.6 62.1 61.8 53.7 33.8 32.8 21.2 53.7
Qwen2.5-7B
RAG with CoT BM25 49.9 47.8 36.0 29.8 36.4 27.4 14.3 16.8 8.7 31.3
E5 52.4 49.4 37.6 33.5 39.1 30.1 16.9 18.8 9.9 34.3
BGE 55.1 51.7 39.6 35.3 40.5 32.0 18.3 20.9 11.3 36.2
ReAct(Tool Call) BM25 43.4 36.0 29.4 28.7 25.1 23.0 11.3 7.6 5.5 27.8
E5 27.9 23.1 21.1 38.3 31.7 26.1 11.7 8.0 5.8 26.0
BGE 37.7 31.4 26.0 27.4 22.6 20.6 11.9 8.2 5.8 25.7
IRCoT BM25 48.5 42.4 36.1 33.8 31.7 29.6 12.8 10.7 9.1 31.7
E5 48.4 42.1 35.7 35.8 33.5 31.1 13.5 11.2 9.4 32.6
BGE 40.5 35.3 30.0 20.1 18.9 18.2 9.2 7.7 6.5 23.3
R3-RAG-CS BM25 60.4 55.5 43.5 49.8 50.4 43.5 25.1 25.3 16.2 45.1
E5 63.3 57.6 45.2 53.1 53.5 46.1 31.3 31.9 21.9 49.2
BGE 64.1 58.7 46.1 55.2 55.2 48.1 31.1 31.1 21.1 50.1
R3-RAG BM25 63.8 58.2 44.9 59.6 61.1 52.8 29.2 30.0 17.6 50.9
E5 65.5 59.7 46.4 62.3 62.7 54.2 33.6 34.0 21.4 53.8
BGE 66.4 60.6 46.8 63.0 63.4 55.2 34.8 34.3 21.7 54.7
Table 4: Comprehensive performance comparison of different retrieval methods (BM25, E5-base-v2, BGE-Large) across all models on HotpotQA, 2WikiMultiHopQA, and MuSiQue datasets. Results include accuracy (ACC), F1, and exact match (EM) metrics.
We present a comprehensive evaluation of datasets and metrics in Table 4. The results show that our model consistently outperforms all baselines across all metrics on the three datasets. Moreover, R3RAG consistently surpasses R3-RAG-CS, further demonstrating the transferability of R3-RAG’s learned reasoning and retrieval strategies across external retrieval environments.
Interestingly, among the baseline methods, both models struggle to effectively utilize the stronger BGE retriever, often performing better with the theoretically weaker BM25 and E5 retrievers. We propose two explanations for this counter-intuitive phenomenon. First, different models have varying adaptation capabilities to retriever characteristics, meaning that stronger general-purpose retrievers don’t always provide better document recall for all models and baseline query formulations. Second, our analysis of retrieval queries and returned documents reveals that baseline methods typically generate broad queries containing multiple knowledge points, expecting retrievers to return comprehensive information in a single operation. However, these knowledge points often have low semantic similarity to each other (e.g., different entity names). BM25 and E5, especially BM25, tend to prioritize documents with high relevance to individual knowledge points, while BGE attempts to optimize for average relevance across all query components, potentially resulting in lower relevance for each specific knowledge point and consequently fewer critical documents compared to BM25.
Appendix G Potential Risk
Previous works have shown LLMs can have various kinds of biasBender et al. (2021). Since our method uses training data distilled from DeepSeek-v3 during the cold-start phase, it can also inherit such biases.
Appendix H Details of Research Artifacts and Licenses
This research utilizes several publicly available datasets, pre-trained models, and frameworks. We provide comprehensive details on their licenses and our usage to ensure transparency and compliance with ethical research standards.
H.1 Datasets and Licenses
We utilized three multi-hop QA datasets:
  • HotpotQA Yang et al. (2018) - Distributed under the CC BY-SA 4.0 license
  • 2WikiMultiHopQA Ho et al. (2020) - Distributed under the Apache-2.0 license
  • MuSiQue Trivedi et al. (2022) - Distributed under the CC BY 4.0 license
Our usage of these datasets adheres strictly to their intended research purposes as specified by their creators. All datasets were used in their original form for academic research purposes only, and no modifications were made to their licenses. The datasets contain questions derived from Wikipedia content and are designed specifically for multi-hop question answering research.
For the artifacts we created in this study (fine-tuned R3-RAG-CS and R3-RAG models), we maintain the same licensing terms as their base models and specify that they are intended for academic research purposes only. These derivative models are compatible with the original access conditions of their base models and datasets, and should not be used outside of research contexts. Our modifications were limited to parameter fine-tuning and did not alter the fundamental architecture or intended use cases of the original models.
H.2 Models, Frameworks, and Their Licenses
We employed the following models and frameworks:
Language Models:
  • Llama-3.1-8B Grattafiori et al. (2024) - Available under the Llama 3.1 Community License
  • Qwen2.5-7B Team (2024) - Available under the Apache-2.0 license
Retriever:
  • E5-base-v2 Wang et al. (2022) - Available under the MIT license
  • BGE-large-en-v1.5 Xiao et al. (2023) - Available under the MIT license
  • BM25S Lù (2024) - Available under the MIT license
Frameworks:
  • LLaMA-Factory Zheng et al. (2024) - Available under the Apache-2.0 license
  • OpenRLHF Hu et al. (2024) - Available under the Apache-2.0 license
  • FlashRAG Jin et al. (2024) - Available under the MIT license
  • FAISS Johnson et al. (2019) - Available under the MIT license
Our use of these artifacts is consistent with their intended purposes. The language models and retrieval models were used for natural language processing and information retrieval tasks within a research context. The frameworks were employed as development tools to implement our research methodology. No commercial applications were developed using these resources. All model adaptations created during this research maintain compatibility with the original license terms of their respective base models.
def prompt_question_init(question):
prompt = f’’
# Role
You are an expert in large language models and knowledge retrieval, with extensive expertise in problem decomposition.
# Answer Format:
The problem analysis: [Analyze information based on the question.]
**Either:**
The retrieval query: [If information is insufficient, generate one query based on the analysis.]
**OR:**
The final answer: [If information is sufficient, provide the final answer based on the analysis.]
# Instructions:
1. **The Problem Analysis**:
- **Decision Making**:
- **Sufficient Information**: If the available information is adequate and no more query need to retrieve, use your parameter knowledge and reasoning capability to solve it and get the answer. Finaly, output the answer in the "The final answer:" section.
- **Insufficient Information**: If the information is lacking, identify what information is missing:
- **Determine Next Retrieval**: Consider what content should be retrieved next and decomposite the query into simpler retrieval queries.
2. **Response Formatting**:
- **Start with Analysis**: Always begin with "The problem analysis:" followed by the detailed analysis. Only one analysis is permitted; therefore, include all your analytical content within this section.
- **Choose Appropriate Section**: Based on the decision:
- Include "The retrieval query:" if more information is needed.
- Include "The final answer:" if sufficient information is available.
- **Generate a Single Query**: Based on the analysis, formulate one targeted query and place it in the "The retrieval query:" section. Ensure that only one query is generated, focusing on the first retrieval question needed.
- **Language Consistency**: Provide the entire response in English, regardless of the input language.
# Key Competencies
1. **Decision Making**:
- **Assessment**: Evaluate whether it is feasible to resolve the entire problem based on all preceding steps.
- **Action in Uncertainty**: If uncertain, initiate an attempt to solve the problem.
- **Failure Handling**: If the solution process encounters a failure, terminate the attempt and analyze which information should be retrieved next.
2. **Decomposition of Retrieval Content**:
- **Problem Breakdown**: When analysis requires solving a problem, decompose complex questions into smaller, actionable sub-questions (atomic problems).
- **Step-by-Step Resolution**: Address these sub-questions sequentially and establish a logical sequence to answer the original question.
- **Identifying Broad Problems**: If the analysis reveals that the information retrieved in the last step does not align with the query, this indicates that the problem is too broad. In such cases:
- **Suggestion**: Propose a method to decompose the problem into smaller sub-questions.
- **Implementation**: Attempt to break down the original query into simpler, more manageable questions.
- **Methods of Decomposition**: Complex questions can be decomposed into **two fundamental forms**:
- **Sequential Decomposition**:
- **Definition**: Break the problem into a series of dependent steps, where each step relies on the result of the previous one.
- **Parallel Decomposition**:
- **Definition**: Split the problem into multiple independent sub-problems, solve each separately, and then combine the results.
#### Current Input:
The question: {question}
#### Output:
return prompt.strip()
Figure 6: The initial prompt template used for the first step of trajectory generation.
def prompt_question(steps_formatted, question):
return f’’
# Role
You are an expert in LLMs and RAG, with extensive expertise in decomposition.
# Answer Format:
The problem analysis: [Analyze information based on previous steps and question]
**Either:**
The retrieval query: [If information is insufficient, generate one query]
**OR:**
The final answer: [If information is sufficient, provide the final answer]
# Instructions:
1. **The Problem Analysis**:
- **Review Previous Steps**: Examine all prior steps, including analyses, queries, and documents.
- **Decision Making**:
- **Sufficient Information**: If the information is adequate, analyze it to solve the problem and output the answer in the "The final answer:" section.
- **Insufficient Information**: If the information is lacking:
- **Evaluate Last Step**: Analyze the documents obtained in the last step to determine if they are relevant to the query or the problem at hand.
- **Utilizing parameter knowledge**: If the retrieval results of the query are not relevant, check if the model’s own parameter knowledge can response the query.
- **Assess Previous Steps**: Identify what information is missing from the overall previous steps.
- **Determine Next Retrieval**: Consider what content should be retrieved next and decomposite the query into simpler retrieval queries.
2. **Response Formatting**:
- **Start with Analysis**: Always begin with "The problem analysis:" followed by the detailed analysis. Only one analysis is permitted; therefore, include all your analytical content within this section.
- **Choose Appropriate Section**: Based on the decision:
- Include "The retrieval query:" if more information is needed.
- Include "The final answer:" if sufficient information is available.
- **Generate a Single Query**: Based on the analysis, formulate one targeted query and place it in the "The retrieval query:" section. Ensure that only one query is generated, focusing on the first retrieval question needed.
- **Language Consistency**: Provide the entire response in English, regardless of the input language.
# Key Competencies
1. **Decision Making**:
- **Assessment**: Evaluate whether it is feasible to resolve the entire problem based on all preceding steps.
- **Action in Uncertainty**: If uncertain, attempt to solve the problem.
- **Failure Handling**: If the solution process encounters a failure, terminate the attempt and analyze which information should be retrieved next.
2. **Decomposition of Retrieval Content**:
- **Problem Breakdown**: When analysis requires solving a problem, decompose complex questions into smaller, actionable sub-questions (atomic problems).
- **Step-by-Step Resolution**: Address these sub-questions sequentially and establish a logical sequence to answer the original question.
- **Identifying Broad Problems**: If the analysis reveals that the information retrieved in the last step does not align with the query, this indicates that the problem is too broad. In such cases:
- **Suggestion**: Propose a method to decompose the problem into sub-questions.
- **Implementation**: Attempt to break down the original query into simpler, more manageable questions.
- **Methods of Decomposition**: Complex questions can be decomposed into **two fundamental forms**:
- **Sequential Decomposition**:
- **Definition**: Break the problem into a series of dependent steps, where each step relies on the result of the previous one.
- **Parallel Decomposition**:
- **Definition**: Split the problem into multiple independent sub-problems, solve each separately, and then combine the results.
#### Current Input:
The question: {question}. The Previous Steps:{steps_formatted.strip()}
#### Output:
Figure 7: The prompt template used for subsequent steps in trajectory generation.